# Learning Lens Blur Fields

### v1

@@ VI Concluding Remarks
We are excited about several avenues of future work. For example, given that we find significant differences between lens blur fields from devices of the same make and model, how distinct are these differences across a wider cross-section of device types and instances, and what can we learn about devices from these aberrations? Additionally, we wish to explore how the lens blur field can be extended to even higher dimensions (e.g., light fields, hyperspectral data, polarization imaging, etc.). Our current approach aims to balance competing goals of reconstructing high-dimensional lens blur fields and achieving an easy-to-use, lightweight method, and so was primarily demonstrated for 5D calibration. However, we are optimistic that the approach can generalize to higher dimensions.
@@
Figure 6: (A) Comparison between calibrated PSFs from the wide cameras of two iPhone 14 Pro devices. We estimate PSFs using four separate sets of calibration measurements captured independently for each device, for each repeatability investigation. Visualizing slices of the mean estimated 5D PSF function for each device (top) reveals broad agreement in the PSF structure; however, significant variations can be seen upon close inspection (insets, bottom). While imperfections in the calibration and capture process result in differences between PSF estimates for the same device (see standard deviation of PSFs, bottom left and right), this effect is generally weaker than PSF differences between devices (bottom center, shown as root-mean-square deviation). (B) Comparison between static and dynamic experiments for a Canon EF 50mm f/1.4 lens. We capture five separate sets of measurements for each repeatbility experiment (with and without remount) and visualize the mean and standard deviations for each experiment with the same color scale. We observe little differences between experiment, indicating that remounting the lens does not alter the blur field of the lens.
@@
Figure 8: Comparison of captures and synthetically-generated renderes with a learned blur field from the iPhone 12 Pro wide lens. A resolution chart (Edmund Optics I3A/ISO 12233 Digital Camera Resolution Test Chart, #58-940) is positioned at a distance of 36 cm from the iPhone 12 Pro, and the blur field is visualized in  Figure 4. The chart is scanned at a resolution of 600 DPI and subsequently blurred at the different focus settings indicated, using the previously-captured blur field. The 2D PSF at the center of each region is used for blurring and shown as insets. Specifically, we examine a zoomed-in region of the chart containing a Siemens star. It is important to note that while some minor demosaicing artifacts and noise are present in the captured image, there is overall agreement between the captured and estimated images across the entire focus range.

### v2

@@ VI Concluding Remarks
We are excited about several avenues of future work. For example, given that we find significant differences between lens blur fields from devices of the same make and model, how distinct are these differences across a wider cross-section of device types and instances? What can we learn about the devices from these aberrations? Additionally, we wish to explore how the lens blur field can be extended to even higher dimensions (e.g., light fields, hyperspectral data, polarization imaging, etc.).
@@
Figure 6: (A) Comparison between calibrated PSFs from the wide cameras of two iPhone 14 Pro devices. We estimate PSFs using four separate sets of calibration measurements captured independently for each device, for each repeatability investigation. Visualizing slices of the mean estimated 5D PSF function for each device (top) reveals broad agreement in the PSF structure; however, significant variations can be seen upon close inspection (insets, bottom). While imperfections in the calibration and capture process result in differences between PSF estimates for the same device (see standard deviation of PSFs, bottom left and right), this effect is generally weaker than PSF differences between devices (bottom center, shown as root-mean-square deviation), suggesting that each device has a specific blur signature. (B) Comparison between static and dynamic experiments for a Canon EF 50mm f/1.4 lens. We capture five separate sets of measurements for each repeatbility experiment (with and without remount) and visualize the mean and standard deviations for each experiment with the same color scale. We observe little differences between experiment, indicating that remounting the lens does not alter the blur field of the lens.
@@
Figure 8: (A) Comparison of real captures and synthetically generated renders using a learned blur field from the iPhone 12 Pro wide lens, and (B) restoration results from these captures via non-blind deconvolution using our blur field PSFs versus a Seidel model fitted to our PSFs. A resolution chart (Edmund Optics I3A/ISO 12233 Digital Camera Resolution Test Chart, #58-940) is positioned at a distance of 36 cm from an iPhone 12 Pro. The chart is scanned at a resolution of 600 DPI and subsequently blurred at the different focus settings indicated, using the previously-captured blur field (visualized in  Fig. 4). The 2D PSF at the center of each region is used for blurring and shown as insets. There is overall agreement between the captured and estimated images across the entire focus range. In the restoration results, we perform deconvolution using the proposed method. Our forward model convolved the ground truth estimate with PSFs to create a synthetic image, which is then compared to the actual capture to compute the loss. The blurry capture is boxed in orange, and the PSFs from blur fields (boxed in blue) and Seidel are shown for each color channel. While the Seidel PSFs are comparable in size to the blur field PSFs, they do not capture real-life effects as effectively. The lens blur field preserves high-frequency details in the Siemens Star arms better than the Seidel PSF, as seen in the yellow-boxed region. The synthetic blurry image generated with blur field PSFs preserves more detail and more closely resembles the captured blurry image than that generated with the Seidel PSFs.
Acknowledgment
DBL and KNK acknowledge support of NSERC under the RGPIN, RTI, and USRA programs. DBL also acknowledges support from the Canada Foundation for Innovation and the Ontario Research Fund.